# Diffusive hydrodynamics from long-range correlations

### v2

@@ Abstract
We illustrate the results for the paradigmatic hard-rod model. It consists on a one dimensional gas of $N$ billiard balls of diameter $a$ and equal unit mass $m=1$. The rods have ordered positions $\{x^{\rm micro}_{i}\}_{i=1}^{N}$ and move freely with velocities $\{\theta_{i}\}_{i=1}^{N}$. The rods collide elastically whenever $x^{\rm micro}_{i+1}-x^{\rm micro}_{i}=a$, simply exchanging their velocities. This model is integrable, since all the initial velocities are preserved by the dynamics Spohn (1991a); Doyon and Spohn (2017). In particular, the charge densities are explicitly defined as $\mathcal{q}_{i}(x,t)=\int{\rm d}{\theta}\,\theta^{i}\rho(x,\theta,t)$, where $\rho(x,\theta,t)$ is the density of particles in space and rapidity .
We focus on the dynamics of the hard-rod gas with initial rapidity density distribution $\rho(x,\theta,t=0)=\exp[-(\theta+\tanh(x))^{2}/2\sigma]/\sigma\sqrt{2\pi}$, where $x=\ell x^{\rm micro}$ and $t=\ell t^{\rm micro}$ are macroscopic time and space coordinates. This state represents a smoothened version of the sharp partition protocol, where the average particle velocity changes from $1$ at $x=-\infty$ to $-1$ at $x=\infty$. Note that the density of particles in the initial state is homogeneous $\mathcal{q}_{0}(x,0)=1$. In a local GGE state with constant particle density $\mathcal{q}_{0}$, the hard rods’ positions do not depend on the momenta and are given by a Poisson point process in the volume-excluded coordinates $\hat{x}_{i}=x_{i}-ai$ Boldrighini and Suhov (1997). Thus, we can generate initial particle positions as $x_{i+1}^{\rm micro}=x_{i}^{\rm micro}+a+(1/\mathcal{q}_{0}-a)\xi_{i}$, where $\xi_{i}$ are i.i.d. standard exponentially distributed variable. The rapidities are then chosen randomly according the to the local rapidity distribution. The dynamics of the momentum distribution $\mathcal{q}_{1}(x,t)$ is shown in Fig. 1 c).
@@
Discussion.—  We introduced a new theory for diffusion in systems with linearly degenerate ballistic modes, particularly applicable to integrable models. Unlike the NS equation, which describes the averages of charge-densities, our theory consists of two coupled equations for both the averages and correlations of charge-densities. This approach reveals that the Kubo-formula based diffusion fails at late times due to the violation of the local equilibrium assumption by correlations. We numerically verified our theory in the hard rods case and found it consistent with independent derivations Hübner et al. .
@@ Acknowledgements.
0.1 Microscopic correlations in the hard rod gas
In this section, we show numerical results for the microscopic correlation functions in a hard rods gas. Let us consider a homogeneous hard rods gas, with rods’ length $a$ and particle density $\rho(\theta)=\bar{\rho}h(\theta)$, with $\int{\rm d}{\theta}h(\theta)=1$ and $\theta$ the particles’ velocities. The microscopic correlation function for such a system is given by Flicker (1968)
$S^{\rm micro}_{\theta,\theta^{\prime}}(x,x^{\prime})=\delta(x-x^{\prime})\delta(\theta-\theta^{\prime})\bar{\rho}h(\theta)+(\mathfrak{n}^{(2)}(x-x^{\prime})-\bar{\rho}^{2})h(\theta)h(\theta^{\prime}).$ (S1)
$\mathfrak{n}^{(2)}(x-x^{\prime})=\frac{\bar{\rho}^{2}}{1-\bar{\rho}a}\sum_{k=1}^{\infty}\frac{1}{(k-1)!}\left(\frac{|x-x^{\prime}|-ak}{\bar{\rho}^{-1}-a}\right)^{k-1}\exp\bigg{[}-\left(\frac{|x-x^{\prime}|-ak}{\bar{\rho}^{-1}-a}\right)\bigg{]}\Theta_{H}(|x-x^{\prime}|-ka)\,,$ (S2)
where $\Theta_{H}$ is the Heaviside step function. Let us now consider a non-uniform gas with $\ell$ the typical scale of spatial variations. Its correlations will be given by $\langle\rho(x,\theta)\rho(x,\theta^{\prime})\rangle^{c}=S^{\rm micro}_{\theta,\theta^{\prime}}(x,x^{\prime})+\mathcal{O}(\ell^{-1})$, where $S^{\rm micro}_{\theta,\theta^{\prime}}(x,x^{\prime})$ is defined by taking $\bar{\rho}\to\bar{\rho}(x,t)$ with $\bar{\rho}(x,t)\equiv\int{\rm d}{\theta}\rho(x,\theta,t)$ and $h(\theta)\to h_{x}(\theta,t)\equiv\rho(x,\theta,t)/\bar{\rho}(x,t)$. We also stress that here $(x,x^{\prime})$ are microscopic spatial coordinates. In a similar way, we can compute the miscroscopic correlations in terms of filling functions $\langle n(x,\theta)n(x,\theta)\rangle^{c}=S^{n,{\rm micro}}_{\theta,\theta^{\prime}}+\mathcal{O}(\ell^{-1})$ where $n(x,\theta)=\rho(x,\theta)/\rho^{\rm tot}(x)$ and with $\rho^{\rm tot}(x)=(1-a\bar{\rho}(x))$,
$S^{n,{\rm micro}}_{\theta,\theta^{\prime}}=(1-a\bar{\rho})^{-2}\bigg{(}\delta(x-x^{\prime})\Big{[}\delta(\theta-\theta^{\prime})\rho(\theta)-a(a\bar{\rho}-2)\frac{\rho(\theta)\rho(\theta^{\prime})}{(1-a\bar{\rho})^{2}}\Big{]}+\Big{(}\frac{\mathfrak{n}^{(2)}(x-x^{\prime})}{\bar{\rho}^{2}}-1\Big{)}\frac{\rho(\theta)\rho(\theta^{\prime})}{(1-a\bar{\rho})^{2}}\bigg{)}\,.$ (S3)
We compare the numerical results for the microscopic correlations functions of a time evolving hard rod gas with the prediction of Eq. (S1). More precisely, we consider a gas of hard rods with only two modes, $\theta_{+}=+1$ and $\theta_{-}=-1$, evolving from the initial state $\rho_{\pm}(x,t=0)=1\mp{\rm Erf}(x/\ell)/2$. In the initial state also $q_{0}(x,t=0)=\rho_{+}(x,t=0)+\rho_{-}(x,t=0)=1$ and $q_{1}(x,t=0)=\rho_{+}(x,t=0)-\rho_{-}(x,t=0)=-{\rm Erf}(x/\ell)$.
Figure S1: (a) Plot for the microscopic correlations in a Hard rods gas, with rods length $a=0.3$ and scale $\ell=200$. Respectively, from left to right, we plot $\langle q_{0}(x,t)q_{1}(y,t)\rangle^{c}_{\rm reg}-\delta(x-y)\langle q_{1}(x,t)\rangle$ at the points $y=\ell/2$, $t=\ell/10$ and $\ell=200$. The Hard rods data (black line) is compared with prediction from Eq. (S1) (red line). The Hard rod data are averaged over $10^{7}$ initial states. Eq. (S1) has been evaluated truncating the summation at $k=50$, such that the truncation error is $<10^{-16}$. The agreement between the prediction and the numerical data is excellent, up to the Hard rods monte carlo noise. We stress that discrepancies induced by long-range correlations are expected to be order $\mathcal{O}(\ell^{-1})$. (b) Plot of $\langle q_{1}(x,t)\rangle$ at a time $t=\ell/10$ from Hard rods numerics. The figure (c) shows the microscopic correlations $\langle q_{0}(x,t)q_{1}(y,t)\rangle^{c}_{\rm reg}-\delta(x-y)\langle q_{1}(x,t)\rangle$ at $y=0$ and $t=10/\ell$. At this point, the leading contribution in $\ell$ is expected to be vanishing, having $\langle q_{0}(x=0,t=\ell/10)\rangle\simeq 0$ (as shown in box (b)), and since $\langle q_{0}(x,t)q_{1}(y,t)\rangle^{c}_{\rm reg}\rangle\propto\langle q_{1}(x,t)\rangle+\mathcal{O}(\ell^{-1})$. Hence, this plots shows the microscopic structure of the discontinuity in the two-point correlation. We can conclude that the ’jump’ is developed in a length scale $\sim a$.
In Fig. S1(a) we show the Hard rods numerical results for $\langle q_{0}(x,t)q_{1}(y,t)\rangle^{c}-\delta(x-y)\langle q_{1}(x,t)\rangle$ at the points $y=\ell/2$ and $t=\ell/10$ for $\ell=200$. Comparing it with the theoretical predictions fro Eq. (S1), we observe excellent agreement. We also stress that all the discrepancies induced by the long-range correlations are at order $\mathcal{O}(\ell^{-1})$, hence they are not visible in Fig. S1 (a). More precisely, we used the relation
$\langle q_{0}(x,t)q_{1}(y,t)\rangle^{c}=\sum_{i,j=1}^{2}S^{\rm micro}_{\theta_{i},\theta_{j}}(x,y,t)\theta_{i}\theta_{j}+\mathcal{O}(\ell^{-1})\quad,\mbox{with:}\quad\sum_{i,j=1}^{2}S^{\rm micro}_{\theta_{i},\theta_{j}}(x,y,t)\theta_{i}\theta_{j}\propto\langle q_{1}(y,t)\rangle$ (S4)
where we used $\theta_{i}=[-1,+1]$ and where $S^{\rm micro}$ is defined in Eq. (S1). But, as it is possible to observe from Fig. S1 (b), $\langle q_{1}(y,t)\rangle$ is vanishing at $t=\ell/10$ and $y=0$. Thus, we expect the leading contribution to be vanishing at this point. This permits to isolate the long-range contribution to the correlations at microscopic scale and to observe the microscopic structure of the discontinuity introduced in the main text. In particular, in Fig. S1 (c), we show the two point function $\langle q_{0}(x,t)q_{1}(y,t)\rangle^{c}-\delta(x-y)\langle q_{1}(x,t)\rangle$ at $t=\ell/10$ and $y=0$. From this picture, we can observe that the discontinuity predicted at hydrodynamic scale, is developed in the gas at scales $\sim a$.
@@ Appendix 1 Dynamics of two-point correlations
$\frac{\delta\mathcal{j}_{i}(\mathcal{q}_{\cdot})}{{\delta\mathcal{q}_{r}\delta\mathcal{q}_{k}}}\,E^{\rm sym}_{rk}$ $=\frac{\delta\mathcal{j}_{i}(\mathcal{q}_{\cdot})}{{\delta\mathcal{q}_{r}\delta\mathcal{q}_{k}}}\,\frac{K_{rk}(x,x;t)}{2}\,\frac{\operatorname{sgn}(u_{r}(x^{-},t)-u_{k}(x^{+},t))+\operatorname{sgn}(u_{r}(x^{+},t)-u_{k}(x^{-},t))}{2}$ (S22)
$=\frac{\delta\mathcal{j}_{i}(\mathcal{q}_{\cdot})}{{\delta\mathcal{q}_{r}\delta\mathcal{q}_{k}}}\,\frac{K_{rk}(x,x;t)}{2}\,\operatorname{sgn}(v^{\rm eff}_{k}-v^{\rm eff}_{r})$ (S23)
$=-\tfrac{1}{2}\langle q_{a},q_{b},j_{i}^{-}\rangle(x,t)C^{ar}C^{bk}(R^{-1})_{r}^{~{}c}(R^{-1})_{k}^{~{}d}\,\frac{\langle n_{c},n_{d},j^{-}_{k}\rangle(x,t)\,C^{kl}\partial_{x}q_{l}(x,t)}{v_{r}^{\rm eff}(x,t)-v_{k}^{\rm eff}(x,t)}\,\operatorname{sgn}(v^{\rm eff}_{k}-v^{\rm eff}_{r})$ (S24)
$=\frac{\langle n_{a},n_{b},j_{i}^{-}\rangle(x,t)\langle n_{r},n_{k},j^{-}_{k}\rangle(x,t)\,C^{kl}\partial_{x}q_{l}(x,t)}{2|v_{r}^{\rm eff}(x,t)-v_{k}^{\rm eff}(x,t)|}.$ (S25)
@@ Appendix 3 Long-Range equilibrium states
$\rho_{\rm LR}\propto$ $\exp\Big{[}-\int{\rm d}x\,\beta^{i}(x)q_{i}(x)+\int_{x\neq y}{\rm d}x{\rm d}y\,\frac{\gamma^{ij}(x,y)}{2}\delta{q}_{i}(x)\delta{q}_{j}(y)\Big{]},$ (S40)
@@
[…]
@@
×$\times\int{\rm d}{p^{\prime\prime}}{\rm d}{q^{\prime\prime}}(\delta(p^{\prime}-p^{\prime\prime})+a\tfrac{\rho_{\mathrm{E}}(x^{\prime},p^{\prime})}{1^{\mathrm{dr}}})(\delta(q^{\prime}-q^{\prime\prime})+\tfrac{\rho_{\mathrm{E}}(y^{\prime},q^{\prime})}{1^{\mathrm{dr}}})$ (S70)
×$\times\Big{(}\int_{-\infty}^{x^{\prime}\wedge y^{\prime}}{\rm d}{z}\rho(z,p^{\prime\prime})\delta(p^{\prime\prime}-q^{\prime\prime})+\rho_{\mathrm{E}}(z,p^{\prime\prime})\rho_{\mathrm{E}}(z,q^{\prime\prime})(-2a+a^{2}\bar{\rho})\Big{)}\Big{|}_{x^{\prime}=\hat{X}^{-1}(0,\hat{X}(t,x)-p^{\prime}t),y^{\prime}=\hat{X}^{-1}(0,\hat{X}(t,y)-q^{\prime}t)}\Big{]},$ (S71)
with $\hat{X}(t,x)=x-a\int_{-\infty}^{x}{\rm d}{y}{\rm d}{p}\rho_{\mathrm{E}}(t,y,p)$ and $\hat{X}^{-1}(t,\hat{x})$ is its inverse in $x$. These formulas are evaluated numerically.

### v3

@@ Abstract
We illustrate the results for the paradigmatic hard-rod model. It consists of a one-dimensional gas of $N$ billiard balls of diameter $a$ and unit mass $m=1$. The rods have ordered positions $\{x^{\rm micro}_{i}\}_{i=1}^{N}$ and move freely with velocities $\{\theta_{i}\}_{i=1}^{N}$. The rods collide elastically whenever $x^{\rm micro}_{i+1}-x^{\rm micro}_{i}=a$, simply exchanging their velocities. This model is integrable, since all the initial velocities are preserved by the dynamics Spohn (1991a); Doyon and Spohn (2017). In particular, the charge densities are explicitly defined as $\mathcal{q}_{i}(x,t)=\int{\rm d}{\theta}\,\theta^{i}\rho(x,\theta,t)$, where $\rho(x,\theta,t)$ is the density of particles in space and rapidity .
We focus on the dynamics of the hard-rod gas with initial rapidity density distribution $\rho(x,\theta,t=0)=\exp[-(\theta+\tanh(x))^{2}/2\sigma]/\sigma\sqrt{2\pi}$, where $x=x^{\rm micro}/\ell$ and $t=t^{\rm micro}/\ell$ are macroscopic time and space coordinates. This state represents a smoothened version of the sharp partition protocol, where the average particle velocity changes from $1$ at $x=$-$\infty$ to 1-1 at $x=\infty$. Note that the density of particles in the initial state is homogeneous $\mathcal{q}_{0}(x,0)=1$. In a local GGE state with constant particle density $\mathcal{q}_{0}$, the hard rods’ positions do not depend on the momenta and are given by a Poisson point process in the volume-excluded coordinates $\hat{x}_{i}=x_{i}-ai$ Boldrighini and Suhov (1997). Thus, we can generate initial particle positions as $x_{i+1}^{\rm micro}=x_{i}^{\rm micro}+a+(1/\mathcal{q}_{0}-a)\xi_{i}$, where $\xi_{i}$ are i.i.d. standard exponentially distributed variable. The rapidities are then chosen randomly according the to the local rapidity distribution. The dynamics of the momentum distribution $\mathcal{q}_{1}(x,t)$ is shown in Fig. 1 c).
@@
Discussion.—  We introduced a new theory for diffusion in systems with linearly degenerate ballistic modes, particularly applicable to integrable models. Unlike the NS equation, which describes the averages of charge-densities, our theory consists of two coupled equations for both the averages and correlations of charge-densities. This approach reveals that the Kubo-formula based diffusion fails at late times due to the violation of the local equilibrium assumption by correlations. We numerically verified our theory in the hard-rod gas and found it consistent with independent derivations Hübner et al. . Note that in integrable systems, the Onsager matrix was evaluated exactly for the first time in De Nardis et al. (2018). Our linear-response results agree with this Onsager matrix, but the claim made there for the full diffusive-scale hydrodynamic equation, which was based on Kubo-like diffusion, is only a good approximation at early times.
@@ Appendix 1 Dynamics of two-point correlations
$\frac{\delta\mathcal{j}_{i}(\mathcal{q}_{\cdot})}{{\delta\mathcal{q}_{r}\delta\mathcal{q}_{k}}}\,E^{\rm sym}_{rk}$ $=\frac{\delta\mathcal{j}_{i}(\mathcal{q}_{\cdot})}{{\delta\mathcal{q}_{r}\delta\mathcal{q}_{k}}}\,K_{rk}(x,x;t)\,\frac{\operatorname{sgn}(u_{r}(x^{-},t)-u_{k}(x^{+},t))+\operatorname{sgn}(u_{r}(x^{+},t)-u_{k}(x^{-},t))}{2}$ (S18)
$=\frac{\delta\mathcal{j}_{i}(\mathcal{q}_{\cdot})}{{\delta\mathcal{q}_{r}\delta\mathcal{q}_{k}}}\,K_{rk}(x,x;t)\,\operatorname{sgn}(v^{\rm eff}_{k}-v^{\rm eff}_{r})$ (S19)
$=-\langle q_{a},q_{b},j_{i}^{-}\rangle(x,t)C^{ar}C^{bk}(R^{-1})_{r}^{~c}(R^{-1})_{k}^{~d}\,\frac{\langle n_{c},n_{d},j^{-}_{k}\rangle(x,t)\,C^{kl}\partial_{x}q_{l}(x,t)}{v_{r}^{\rm eff}(x,t)-v_{k}^{\rm eff}(x,t)}\,\operatorname{sgn}(v^{\rm eff}_{k}-v^{\rm eff}_{r})$ (S20)
$=\frac{\langle n_{a},n_{b},j_{i}^{-}\rangle(x,t)\langle n_{r},n_{k},j^{-}_{k}\rangle(x,t)\,C^{kl}\partial_{x}q_{l}(x,t)}{|v_{r}^{\rm eff}(x,t)-v_{k}^{\rm eff}(x,t)|}=-\mathfrak{D}_{i}^{~l}\partial_{x}q_{l}(x,t).$ (S21)
@@ Appendix 2 Degenerate three-point coupling from linear degeneracy
Remark: Plugging (S22) into (S21) we see that linear degeneracy is required for a finite diffusion matrix $\mathfrak{D}_{i}^{~k}$ and the Onsager matrix $\mathfrak{L}_{ij}$.
@@ Appendix 3 Long-Range equilibrium states
$\rho_{\rm LR}\propto$ $\exp\Big[-\int{\rm d}x\,\beta^{i}(x)q_{i}(x)+\int_{x\neq y}\hskip-8.5359pt{\rm d}x{\rm d}y\,\frac{\gamma^{ij}(x,y)}{2}\delta{q}_{i}(x)\delta{q}_{j}(y)\Big],$ (S36)
@@
$\dot{S}=\int dx\partial\beta^{k}(j_{k}(q_{\cdot})+\frac{\gamma_{s}^{ij}(x,x)}{2}\langle j_{k}^{-},q_{i},q_{j}\rangle)-\int{\rm d}x\gamma_{s}^{ij}(x,x)\langle j_{k}^{-},q_{i},q_{j}\rangle\partial\beta^{k}+\int dx\partial_{x}j_{s}=0$ (S54)
@@
$\times\int{\rm d}{p^{\prime\prime}}{\rm d}{q^{\prime\prime}}(\delta(p^{\prime}-p^{\prime\prime})+a\tfrac{\rho_{\mathrm{E}}(x^{\prime},p^{\prime})}{1^{\mathrm{dr}}})(\delta(q^{\prime}-q^{\prime\prime})+\tfrac{\rho_{\mathrm{E}}(y^{\prime},q^{\prime})}{1^{\mathrm{dr}}})$ (S66)
$\times\Big(\int_{-\infty}^{x^{\prime}\wedge y^{\prime}}{\rm d}{z}\rho(z,p^{\prime\prime})\delta(p^{\prime\prime}-q^{\prime\prime})+\rho_{\mathrm{E}}(z,p^{\prime\prime})\rho_{\mathrm{E}}(z,q^{\prime\prime})(-2a+a^{2}\bar{\rho})\Big)\Big|_{x^{\prime}=\hat{X}^{-1}(0,\hat{X}(t,x)-p^{\prime}t),y^{\prime}=\hat{X}^{-1}(0,\hat{X}(t,y)-q^{\prime}t)}\Big],$ (S67)
[…]